Eliminate the arbitrary constants.


y = x + Aex + Be-2x

1
Expert's answer
2021-09-15T02:57:17-0400

Given


"y=x+Ae^x + Be^{-2x} \quad \cdots (i)"


From eqn (i) above,


"y-x=Ae^x + Be^{-2x} \quad \cdots (ii)"

Since we have two arbitrary constants, we'll differentiate eqn (i) twice.


"y'=1+Ae^x -2Be^{-2x} \quad \cdots (iii)\\
y''=Ae^x +4Be^{-2x} \quad \cdots (iv)"

Adding eqn (iii) and eqn (iv), we have:


"y''+y'=1+2Ae^x +2Be^{-2x}\\
y''+y'=1+2(Ae^x +Be^{-2x}) \quad \cdots (v)"

Put (ii) in (v), we get


"y''+y'=1+2(y-x)\\"

Re-write


"y''+y'-2y=1-2x"